# Systematic Study of Amorphous ABC Heterostructures at the Atomic Scale as a Nonlinear Optical Metamaterial

### v1

@@ II Deposition and Characterization
with $\chi^{(1)}$ as the linear susceptibility of the material, to estimate the values at a wavelength of 1032 nm. However, since a femtosecond laser is used, a delay is introduced between the two pulses from the front and back sides of the sample. An SH pulse is generated on the front side of the sample. When it reaches the back side of the sample, another SH pulse is already generated because the pulse of the fundamental wavelength travels at different group velocity through the substrate compared to the SH pulse. This delay between the two SH pulses when they reach the backside of the sample is called the temporal walk-off time $t_{\text{walk-off}}$. This causes destructive interference to be incomplete, and thus, the depth of the interference fringes is decreased. When the temporal walk-off is not taken into account, the value of $\chi_{\textup{}}^{(2)}$ is overestimated. For our fused silica sample, 1 mm thick, the value was measured to be $t_{\text{walk-off}}=94$ fs. This value corresponds well with previous measurements in the literature and also with the calculation of the delay based on the group velocity of each pulse in the substrate. Therefore, the model was modified by Hermans et al. for the pulsed laser with the temporal walk-off effect as
@@
where $K_{2}$ is the calibration constant, $\Delta t$ is the pulse duration at FWHM. However, the walk-off time is not constant; it changes as the angle of incidence increases, and the optical path through the substrate increases as well. Then, the relation is $t_{\text{walk-off}}={t_{\text{walk-off,0}}}/{\cos(\theta_{\text{glass}})}$, where $t_{\text{walk-off,0}}$ is walk-off time at the normal angle of incidence. The walk-off effect is also visualized in Fig. 3. By substituting Eq. 6 with Eq. 3, we can obtain the desired second-order susceptibility together with the other fitting parameters $\chi_{\textup{s},zzz}^{\textup{ABC}}$, $A_{\textup{s},zx}^{\textup{ABC}}$, $K_{2}$, and $L_{\textup{glass}}$. The thickness of the glass substrate must also be fitted for each measurement, as the position of the fringes in the second harmonic is highly sensitive to variations in substrate thickness. Dividing the surface susceptibility by the total thickness of the heterostructure $\chi_{\textup{bulk}}^{\textup{ABC}}=\chi_{\textup{s}}^{\textup{ABC}}/d$, the bulk second-order susceptibility is determined. This simplified version of the model neglects multiple reflections in the film and substrate while maintaining validity, as previously tested. Therefore, we keep the angle of incidence below $60^{\circ}$ for the analysis.
@@
III Results and discussion
To investigate the nonlinear interface response of ABC heterostructures, multiple samples with a constant total deposition thickness of approximately 60 nm were fabricated, while systematically varying the ABC period thickness $t_{\text{ABC}}$ from 60 nm down to the atomic scale of 0.3 nm. This approach enabled the study of nonlinear responses by altering the density of interfaces per unit thickness and exploring different morphologies across the samples, with the goal of understanding how these variations influence the properties of heterostructures. Each oxide material is represented approximately equally in the period $t_{\text{ABC}}$. The ABC heterostructures consist of around 300–380 atomic layers for a 60 nm thick period and only about two atomic layers for an ABC period thickness of 0.3 nm, assuming a Si–O bond length of 0.15 nm, a Ti–O bond length of 0.20 nm, and an Al–O bond length of 0.16 nm Si-O ; Ti-O ; Al-O .
@@ III Results and discussion
Figure 4: ABF STEM images showing the SiO2/TiO2/Al2O3 nanolaminates on the silicon wafer substrate of: (left) sample S3 with the ABC period thickness $t_{\textup{ABC}}$ of 1.5 nm; TiO2 layers appear darker; (right) sample S2 with $t_{\textup{ABC}}$ of 0.7 nm. The layers are no longer visible. The dark area on the top of both nanolaminates is the Pt/C protection layer used for TEM specimen preparation.
To verify the precise thickness of the layers in each composition, X-ray reflectivity (XRR) was employed. The measurement spectra of the heterostructures are presented in Fig. 5. The Bragg peaks are clearly visible for all samples except for the heterostructure with a period of 0.7 nm (sample S2). As the period of the heterostructure decreases, the first Bragg peak shifts to the higher grazing angle. The XRR analysis shows the existence of separated nanolaminates and the excellent agreement with the targeted ABC period thickness. The analysis also measured the total deposition thickness and provided the average interface width, incorporating both roughness and interlayer diffusion, see Table 2. Sample S2 has an interface width of 1.2 nm, greater than other samples, with a period thickness of $t_{\text{ABC}}=0.7$ nm, suggesting notable layer intermixing and a lack of clearly separated layers.
Figure 5: Results of the X-ray reflectivity measurements proving the existence of the layer separation and verifying the period thickness $t_{\text{ABC}}$. Sample ID from the top: S13, S11, S6, S3, S2.
@@
Sample ID Total thickness $d$ (nm) Period thickness $t_{\textup{ABC}}$ (nm) Layer thickness SiO2 (nm) Layer thickness TiO2 (nm) Layer thickness Al2O3 (nm) Interface width (nm)
S2 57.3 0.7 0.4 0.16 0.15 1.2
@@
Sample ID Deposition cycles SiO2/TiO2/Al2O3 Super cycles $M$ Layer thickness (nm) SiO2/TiO2/Al2O3 Period thickness $t_{\textup{ABC}}$ (nm) Total thickness $d$ (nm) $n_{\text{o}}$ (1032nm)(1032\ $\text${nm}) $n_{\text{e}}$ (1032nm)(1032\ $\text${nm}) $N_{\text{o}}$ (516nm)(516\ $\text${nm}) $N_{\text{e}}$ (516nm)(516\ $\text${nm}) $A_{\text{zx}}$ (pm/V) $\chi_{\text{zzz}}$ (pm/V)
[…]
@@
*: No birefringence was measured, $\dagger$: The effective medium approximation is no longer valid in the wavelength range from 400 nm to 1040 nm based on the ellipsometric evaluation Layer and period thickness $t_{\textup{ABC}}$ is calculated based on the growth rate per cycle, and the total thickness $d$ is based on ellipsometric evaluation. The uncertainty for the refractive index is $\pm 0.01$. Samples close to 60 nm of the total thickness are highlighted in gray.
The dispersion curve of the ordinary and extraordinary refractive index is presented in Fig. 6. The magnitude of birefringence, $\Delta n=n_{\text{e}}-n_{\text{o}}$, depends on the thickness of the ABC period, and completely disappears in samples with nonexistent layer separation. Optical properties are summarized in Table 3.
@@
Figure 6: Ordinary and extraordinary refractive indices SiO2/TiO2/Al2O3 nanolaminates of the sample S3 show the presence of the form birefringence.
@@
Polarization measurements were carried out to verify the symmetry properties of the ABC stacks. Figure 7 shows s- and p-polarization of the SH signal from sample S3 as a function of the polarization angle of the fundamental field $\varphi$ with the angle of incidence $\vartheta=63^{\circ}$. S-polarization is represented by $\varphi=0^{\circ}$ and p-polarization by $\varphi=90^{\circ}$. The polarization measurements correspond with the symmetry group $C_{\infty v}$. The maximum SH signal is observed when both the fundamental and the second harmonic are p-polarized.
@@
Figure 7: Normalized SHG power for s- and p-polarizations as a function of the polarization angle $\varphi$ of the fundamental corresponding with the symmetry group $C_{\infty v}$, measured for the sample S3.
@@
Based on the experimental results for the p-polarized fundamental and the second harmonic in Fig. 8(a), the intensity of the second harmonic increases with decreasing ABC period thickness from 60 nm until 1.5 nm, where a maximum is observed. For periods thinner than 1.5 nm, the second harmonic signal decreases abruptly. As each layer is approximately 0.25 nm thick, intermixing at the interfaces likely degrades layer quality, and incomplete layer closure further reduces the SH signal. This hypothesis is supported by STEM imaging and XRR analysis of sample S2, which show no clear layer separation for an ABC period of $t_{\textup{ABC}}=0.7$ nm. The nonlinear model for the determination of $\chi^{\textup{(2)}}$ is fitted to all SHG data across varying ABC periods, with selected curves shown in Fig. 8(b).
@@
Figure 8: (a) Angle-dependent measurements of the second harmonic frequency from SiO2/TiO2/Al2O3 heterostructures deposited on 1 mm thick fused silica substrate. The total deposited thickness is approximately 60 nm across all samples; only the morphology, i.e., the ABC period thickness $t_{\textup{ABC}}$, varies from sample to sample. Both FF and SH are p-polarized with average power $P=0.5$ W. (b) Fitting of SHG data from selected samples to analyze the second-order susceptibility of the ABC heterostructures.
@@
The same trend is also reported for power dependency measurements. The samples are positioned at the angle of incidence when SHG maxima are observed for p-polarized FF and SHG. In addition, an uncoated substrate of fused silica with natural surface second-order nonlinearity is also included to compare it with the SH intensity across the ABC samples. Verification of second-harmonic generation is achieved by observing its quadratic power dependency, and excellent agreement is found by achieving a slope line of $2.026(9)$ for sample S3, Fig. 9. The second harmonic exhibits a quadratic relationship with the ABC super cycles $M$, similar to the classical bulk crystals, Fig. 10. For each ABC period, second-order susceptibility is determined with the results shown in Fig. 11(a). The trend suggests an enhancement of both $\chi_{zzz}$ and $A_{zx}$ with decreasing ABC period thickness and achieving $\chi_{zzz}=2.0\pm 0.2$ pm/V for $t_{\textup{ABC}}=1.5$ nm. For shorter periods, the values of $\chi_{zzz}$ and $A_{zx}$ decrease rapidly. Several thicknesses were analyzed to clarify the properties at different deposited total thicknesses $d$. The analysis shows that the nonlinear susceptibility remains constant for total thicknesses from 60 nm to 233 nm for a period of 1.5 nm. A similar trend is also observed for periods of 3 nm and 6 nm.
@@
Figure 9: Second harmonic power dependence on fundamental power in SiO2/TiO2/Al2O3 heterostructures, measured at the angle of incidence of maximum second harmonic intensity for both p-polarized fundamental and second harmonic. Quadratic power dependence of the second harmonic for sample S3 with period $t_{\textup{ABC}}=1.5$ nm, fitted to the data.
@@
Figure 10: Second harmonic power scales with quadratic dependence on the film thickness and with the number of ABC super cycles $\text{ABC}]\times M$. Both fundamental and SH are p-polarized with average fundamental power $P=0.5$ W. ABC period thickness is $t_{\textup{ABC}}=3.0$ nm for all samples. Measured at the angle of incidence $\vartheta$ where the global maxima occurs.
@@
Our best ABC compositions are compared in Table 4 with other commonly used nonlinear crystals, as well as with previously reported results for ABC type heterostructures in the literature. We achieve higher nonlinearity than previously reported values, and we attribute this improvement to better-optimized PEALD processes. For example, Alloatti et al. ABC-Alloatti report difficulties with the HfO2 process. Moreover, we achieved much thinner and still enclosed layers compared to the previous reports. ABC period thicknesses $t_{\textup{ABC}}$ were 2.1 nm and 2.7 nm, compared to our best sample with 1.5 nm. Material selection may be partially responsible for the higher observed nonlinearity caused by the dielectric contrast between the layers.
@@
Monocrystal SiO2 (Quartz) - 1.53/1.54 $\textit{D}_{3}$ 1064 - $d_{\text{11}}$ = 0.3 lithiumbiobate ; quartz-n
Monocrystal BaB2O4 (BBO) - 1.67/1.54 $\textit{C}_{3v}$ 1064 - $d_{\text{22}}$ = 2.3 lithiumbiobate ; BBO-n
Monocrystal LiNbO3 (LN) - 2.23/2.16 $\textit{C}_{3v}$ 1064 - $d_{\text{33}}$ = -27 lithiumbiobate ; LN-n
ABC Amorphous HfO2/TiO2/Al2O3 2.0/2.1/1.6 - $C_{\infty v}$ 800 2.7 $d_{\text{33}}$ = 0.4 On-determination-SHG
ABC Amorphous In2O3/TiO2/Al2O3 2.2/2.1/1.6 - $C_{\infty v}$ 800 2.1 $d_{\text{33}}$ = 0.6 On-determination-SHG
@@
To further clarify the origin of the effective bulk nonlinearity, two identical ABC samples S3 were placed on top of each other. First, two samples with stacking sequence I) CBA…CBA-Silica/CBA…CBA-Silica and sequence II) CBA…CBA-Silica/Silica-ABC…ABC were investigated, Fig. 12. Both stacking sequences produce almost the same maximum SHG intensity. However, the number of interference fringes is increased for sequence b), and the depth of fringes is reduced. The temporal walk-off effect is more significant because the optical thickness of the substrates between the heterostructures is twice as large. However, when the stacking sequence I) Silica-ABC…ABC/CBA…CBA-Silica is applied, the SHG almost vanishes. This result suggests that the centrosymmetry is restored. However, SH waves must be generated in each ABC and CBA layer; therefore, they are generated in each stack with an opposite phase and destructively interfere. This leads to the conclusion that the two susceptibilities of the ABC and CBA layers must have an opposite sign, for example, $|\chi_{\textup{s}}^{\textup{ABC}}|=|-\chi_{\textup{s}}^{\textup{CBA}}|$. The polarity flipping of the second-order susceptibility by reversing the growing order was also demonstrated in semiconductor heterostructures based on quantum-engineered intersubband transitions, where the polarity is flipped by spatial inversion in the growth sequence chi2flipping , similar to our case. However, the ABC and CBA sequences can differ under real conditions. For example, material A can grow on material B at a different rate and quality than material B on A. It is yet to be determined whether all growth permutations ABC, ACB, BAC, BCA, CAB, and CBA result in equivalent stack characteristics.
@@
Figure 12: (a) Showing each stacking sequence for the two identical samples with total deposited ABC stack thickness of 60 nm and thickness of the fused silica substrate is 1 mm. (b) Showing angle-dependent measurements for three distinct stacking sequences of the ABC sample, both for p-polarized fundamental and second harmonic. Depending on the stacking sequence configuration, the SH response varies significantly.
IV Conclusion

### v4

@@ 2 Deposition and Characterization
with $\chi^{(1)}$ as the linear susceptibility of the material, to estimate the values at a wavelength of 1032 nm. Because a femtosecond laser is used, the group velocity mismatch between the fundamental and its SH introduces a temporal delay between the SH generated at the front and back surfaces of the sample. By the time the SH generated at the front surface reaches the back surface, an additional SH contribution has already formed there. This relative delay between the two SH contributions is known as the temporal walk-off time $t_{\text{walk-off}}$. This causes destructive interference to be incomplete, and thus, the contrast of the interference fringes is decreased. When the temporal walk-off is not taken into account, the value of $\chi_{\textup{}}^{(2)}$ is overestimated. For our fused silica sample, 1 mm thick, the value was measured to be $t_{\text{walk-off}}=94$ fs. This value corresponds also well with the calculation of the delay based on the group velocity of each pulse in the substrate. Therefore, the model was modified by Hermans et al.[15, 14, 17] for the pulsed laser with the temporal walk-off effect as
@@
where $K_{2}$ is the calibration constant, $\Delta t$ is the pulse duration at FWHM. However, the walk-off time is not constant; it changes as the angle of incidence increases, and the optical path through the substrate increases as well. Then, the relation is $t_{\text{walk-off}}={t_{\text{walk-off,0}}}/{\cos(\theta_{\text{glass}})}$, where $t_{\text{walk-off,0}}$ is walk-off time at the normal angle of incidence. The walk-off effect is also visualized in Fig. 3. By substituting Eq. 6 with Eq. 3, we can obtain the desired second-order susceptibility together with the other fitting parameters $\chi_{\textup{s},zzz}^{\textup{ABC}}$, $A_{\textup{s},zx}^{\textup{ABC}}$, $K_{2}$, and $L_{\textup{glass}}$. The thickness of the glass substrate must also be fitted for each measurement, as the position of the fringes in the second harmonic is highly sensitive to variations in substrate thickness. Dividing the surface susceptibility by the total thickness of the heterostructure $\chi_{\textup{bulk}}^{\textup{ABC}}=\chi_{\textup{s}}^{\textup{ABC}}/d$, the bulk second-order susceptibility is determined. This simplified version of the model neglects multiple reflections in the film and substrate while maintaining validity, as previously tested. Therefore, we keep the angle of incidence below $60^{\circ}$ for the analysis. For additional details, see Supplement 1.
@@
3 Results and discussion
Atomic layer deposition (ALD) enables precise control over thin-film thickness via self-limited growth, with growth rates per cycle (GPC) of $0.12$ nn/cycle for SiO2 (A), $0.07$ nn/cycle for TiO2 (B), and $0.15$ nn/cycle for Al2O3 (C). To investigate the nonlinear interface response of ABC heterostructures, multiple samples with a constant total deposition thickness of approximately 60 nm were fabricated, while systematically varying the ABC period thickness $t_{\text{ABC}}$ from 60 nm down to the atomic scale of 0.3 nm and it is easily achieved by ALD by controlling the number of cycles. This approach enabled the study of nonlinear responses by altering the density of interfaces per unit thickness and exploring different morphologies across the samples, with the goal of understanding how these variations influence the properties of heterostructures. Each oxide material is represented approximately equally in the period $t_{\text{ABC}}$. The ABC heterostructures consist of around 300–380 atomic layers for a 60 nm thick period and only about two atomic layers for an ABC period thickness of 0.3 nm, assuming a Si–O bond length of 0.15 nm, a Ti–O bond length of 0.20 nm, and an Al–O bond length of 0.16 nm [24, 21, 38], suggesting ultimate physical limits for ABC layers.
@@ 3 Results and discussion
Figure 4: ADF STEM images showing the SiO2/TiO2/Al2O3 nanolaminates on the silicon wafer substrate of: (left) sample S3 with the ABC period thickness $t_{\textup{ABC}}$ of 1.5 nm; TiO2 layers appear darker; (right) sample S2 with $t_{\textup{ABC}}$ of 0.7 nm. The layers are no longer visible. The dark area on the top of both nanolaminates is the Pt/C protection layer used for TEM specimen preparation.
To verify the precise thickness of the layers in each composition, X-ray reflectivity (XRR) was employed. The measurement spectra of the heterostructures are presented in Fig. 5. The Bragg peaks are clearly visible for all samples except for the heterostructure with a period of 0.7 nm (sample S2), suggesting notable layer intermixing and a lack of clearly separated layers. As the period of the heterostructure decreases, the first Bragg peak shifts to the higher grazing angle. The XRR analysis shows the existence of separated nanolaminates and the excellent agreement with the targeted ABC period thickness. The analysis also provided the total deposition thickness and provided the average interface roughness of all three types of interfaces in the heterostructure, see Table 2.
Figure 5: Results of the X-ray reflectivity measurements proving the existence of the layers and verifying the period thickness $t_{\text{ABC}}$. Sample ID from the top: S13, S11, S6, S3, S2.
@@
Sample ID Total thickness $d$ (nm) Period thickness $t_{\textup{ABC}}$ (nm) Layer thickness SiO2 (nm) Layer thickness TiO2 (nm) Layer thickness Al2O3 (nm) Average interface roughness (nm)
@@
Time-of-flight secondary ion mass spectrometry (TOF-SIMS) depth profiles of the heterostructures are shown in Fig. 6. In sample S3 ($t_{\text{ABC}}=1.5$ nm), periodic modulation is observed only for Si+ ions, indicating distinct SiO2 layers, while the Al+ and Ti+ signals do not show a clear periodicity. In contrast, sample S11 ($t_{\text{ABC}}=5.9$ nm) exhibits periodic modulation of all constituent cations (Si+, Al+, Ti+). The lack of resolved Al2O3 and TiO2 layering in S3 does not rule out separated ABC layers, as it may arise from limited depth resolution due to differential sputtering yields, intermixing, or sample-preparation effects. Independent X-ray reflectivity (XRR) and scanning transmission electron microscopy (STEM) analyzes confirm the presence of separated layers in S3 with the 1.5 nm period.
Figure 6: Depth profiles of positive secondary ion yields measured by TOF-SIMS in the ABC heterostructure: (a) sample S3 and (b) sample S11.
@@
Sample ID Deposition cycles SiO2/TiO2/Al2O3 Super cycles $M$ Layer thickness (nm) SiO2/TiO2/Al2O3 Period thickness $t_{\textup{ABC}}$ (nm) Total thickness $d$ (nm) $n_{\text{o}}$ (1032 nm) $n_{\text{e}}$ (1032 nm) $N_{\text{o}}$ (516 nm) $N_{\text{e}}$ (516 nm) $A_{\text{zx}}$ (pm/V) $\chi_{\text{zzz}}$ (pm/V)
S1* 1:1:1 561 0.12/0.07/0.15 0.3 57 1.76 - 1.80 - 0.29 0.54
S2* 2:3:2 80 0.24/0.21/0.3 0.7 57 1.81 - 1.86 - 0.50 0.96
S3 4:6:4 40 0.48/0.42/0.6 1.5 59 1.77 1.75 1.83 1.81 0.80 2.0
@@
*: No birefringence was observed. $\dagger$: The effective medium approximation is no longer valid in the wavelength range from 400 nm to 1040 nm based on the ellipsometric evaluation. Layer and period thickness $t_{\textup{ABC}}$ is calculated based on the growth rate per cycle, and the total thickness $d$ is based on ellipsometric evaluation. The uncertainty for the refractive index is $\pm 0.01$. Samples with targeted 60 nm of the total thickness are highlighted in gray.
ABC-type heterostructures have birefringent properties [14]. For monocrystals, birefringence can be explained by the anisotropic electrical properties of the molecules that make up the crystals. However, birefringence can also be caused by anisotropy if the material is arranged on an order of magnitude larger than the size of the molecules and this distance is still smaller than the wavelength of light. The ABC-type heterostructure, composed of three materials arranged in thin layers, acts as an effective medium exhibiting form birefringence and functioning as a negative uniaxial crystal. At the same time, the entire structure is amorphous, and the optical axis is perpendicular to the layers [4, 9]. Ellipsometry confirms the presence of the negative birefringence, and the dispersion curve of the ordinary and extraordinary refractive index are presented in Fig. 7. The magnitude of birefringence, $\Delta n=n_{\text{e}}-n_{\text{o}}$, depends on the thickness of the ABC period, achieving up to $\Delta n=0.15$ for sample S12, and completely disappears in samples with nonexistent layers S1 and S2. Optical properties are summarized in Table 3.
@@
Figure 7: Ordinary and extraordinary refractive indices of SiO2/TiO2/Al2O3 nanolaminates of the sample S3 show the presence of the form birefringence.
@@
Nonlinear polarization measurements were conducted to confirm the symmetry properties of ABC stacks. The linear polarization angle of the fundamental field, $\varphi$, was varied, while the second harmonic analyzer was fixed along the s- or p-polarization direction. At $\varphi=0^{\circ}$, the pump light is s-polarized, and at $\varphi=90^{\circ}$, it is p-polarized. Figure 8 displays the s- and p-polarized SH signals from sample S3 as a function of $\varphi$, with an incidence angle of $\vartheta=63^{\circ}$. The polarization data align with the $C_{\infty v}$ symmetry group, with the maximum SH signal observed when both the fundamental and SH fields are p-polarized. The discrepancy between the theoretical and experimental s-polarized SHG data may stem from the strain present in the ABC stack, which alters the nonlinear susceptibility tensor via the photoelastic effect [23].
@@
Figure 8: Normalized SHG power for s- and p-polarizations as a function of the polarization angle $\varphi$ of the fundamental corresponding with the symmetry group $C_{\infty v}$, measured for the sample S3. Dots represent measurements and lines normalized values according to the theory.
@@
Based on the experimental results for the p-polarized fundamental and the second harmonic in Fig. 9(a), the intensity of the second harmonic increases with decreasing ABC period thickness from 60 nm until 1.5 nm, where a maximum is observed. For periods thinner than 1.5 nm, the second harmonic signal decreases abruptly. As each layer is approximately 0.25 nm thick, intermixing at the interfaces likely degrades layer quality, and incomplete layer closure further reduces the SH signal. This hypothesis is supported by STEM imaging and XRR analysis of sample S2, which show no clear layers for an ABC period of $t_{\textup{ABC}}=0.7$ nm. The nonlinear model for the determination of $\chi^{\textup{(2)}}$ is fitted to all SHG data across varying ABC periods, with selected curves shown in Fig. 9(b).
@@
Figure 9: (a) Angle-dependent measurements of the second harmonic from SiO2/TiO2/Al2O3 heterostructures deposited on 1 mm thick fused silica substrate. The total deposited thickness is approximately 60 nm across all samples; only the morphology, i.e., the ABC period thickness $t_{\textup{ABC}}$, varies from sample to sample. Both $F$ with an average pump power $P=0.5$ W and SH are p-polarized. (b) Fitting of SHG data from selected samples to analyze the second-order susceptibility of the ABC heterostructures.
The same trend is also reported for power dependency measurements. The samples are aligned at the incidence angle that yields the maximum SHG signal for p-polarized fundamental light. For comparison, an uncoated substrate of fused silica with surface second-order nonlinearity is also included. Verification of second-harmonic generation is achieved by observing its quadratic power dependency, and excellent agreement is found by achieving a slope line of $2.026(9)$ for sample S3, see Fig. 10.
Several thicknesses were analyzed to clarify the properties at different deposited total thicknesses $d$. The second harmonic exhibits a quadratic relationship with the ABC super cycles $M$, similar to the classical bulk crystals, see Fig. 11. Previously reported studies on ABC composites also identified the bulk-like SHG behavior [2, 7].
@@
For each ABC period, second-order susceptibility is determined by fitting with the results shown in Fig. 12(a). The trend suggests an enhancement of both $\chi_{zzz}$ and $A_{zx}$ with decreasing ABC period thickness and achieving $\chi_{zzz}=2.0\pm 0.2$ pm/V for $t_{\textup{ABC}}=1.5$ nm. For shorter periods, the values of $\chi_{zzz}$ and $A_{zx}$ decrease rapidly. The analysis also shows that the nonlinear susceptibility remains constant for total thicknesses from 60 nm to 233 nm for a period of 1.5 nm. A similar trend is also observed for periods of 3.0 nm and 6.0 nm in Fig. 12(b).
@@
Figure 10: Average second harmonic power dependence on average fundamental power in SiO2/TiO2/Al2O3 heterostructures, measured at the angle of incidence of maximum second harmonic intensity for both p-polarized fundamental and second harmonic. Quadratic power dependence of the second harmonic for sample S3 with $t_{\textup{ABC}}=1.5$ nm, fitted to the data.
@@
Figure 11: Average second harmonic power scales with quadratic dependence on the film thickness and with the number of ABC super cycles $\text{ABC}]\times M$. Both fundamental and SH are p-polarized with average fundamental power $P=0.5$ W. ABC period thickness is $t_{\textup{ABC}}=3.0$ nm for all samples. Measured at the angle of incidence $\vartheta$ where the global maxima occurs.
@@
Our best ABC compositions are compared in Table 4 with other commonly used nonlinear crystals, as well as with previously reported results for ABC-type heterostructures in the literature. We achieve higher nonlinearity than previously reported values, and we attribute this improvement to higher quality thin film heterostructures, higher quality plasma source and ALD processes. For example, Alloatti et al. [2] report difficulties with the HfO2 process and overall structural inhomogeneity of the heterostructure. Moreover, in this systematic study we achieved much thinner, smoother, and still enclosed layers compared to the previous reports [2, 7]. ABC period thicknesses $t_{\textup{ABC}}$ were 2.1 nm and 2.7 nm, compared to our best sample with 1.5 nm. Material selection may be partially responsible for the higher observed nonlinearity caused by the dielectric contrast between the layers.
@@
Monocrystal SiO2 (Quartz) - 1.53/1.54 $\textit{D}_{3}$ 1064 - $d_{\text{11}}$ = 0.3 [31, 10]
Monocrystal BaB2O4 (BBO) - 1.67/1.54 $\textit{C}_{3v}$ 1064 - $d_{\text{22}}$ = 2.3 [31, 8]
Monocrystal LiNbO3 (LN) - 2.23/2.16 $\textit{C}_{3v}$ 1064 - $d_{\text{33}}$ = -27 [31, 43]
ABC Amorphous HfO2/TiO2/Al2O3 2.0/2.1/1.6 - $C_{\infty v}$ 800 2.7 $d_{\text{33}}$ = 0.4 [15]
ABC Amorphous In2O3/TiO2/Al2O3 2.2/2.1/1.6 - $C_{\infty v}$ 800 2.1 $d_{\text{33}}$ = 0.6 [15]
@@
To further clarify the origin of the effective bulk nonlinearity [7], two identical ABC samples S3 were placed on top of each other. First, two samples with stacking sequence I) CBA…CBA-Silica/CBA…CBA-Silica and sequence II) CBA…CBA-Silica/Silica-ABC…ABC were investigated, see Fig. 13. Both stacking sequences produce almost the same maximum SHG power. However, the number of interference fringes is increased for sequence II), and the contrast of the fringes is reduced. In this case, the temporal walk-off effect is more significant because the optical thickness of the substrates between the heterostructures is twice as large. However, when the stacking sequence III) Silica-ABC…ABC/CBA…CBA-Silica is applied, the SHG signal almost vanishes. This result suggests that the overall centrosymmetry of the stack is restored. However, SH waves must be generated in each ABC and CBA layer; therefore, they are generated in each stack with an opposite phase and destructively interfere. This leads to the conclusion that the two susceptibilities of the ABC and CBA layers must have an opposite sign, for example, $\chi_{\textup{s}}^{\textup{ABC}}=-\chi_{\textup{s}}^{\textup{CBA}}$. The polarity flipping of the second-order susceptibility by reversing the growing order was also demonstrated in semiconductor heterostructures based on quantum-engineered intersubband transitions, where the polarity is flipped by spatial inversion in the growth sequence [33], similar to our case. However, the ABC and CBA sequences can differ under real conditions. For example, material A can grow on material B at a different rate and quality than material B on A. It is yet to be determined whether all growth permutations ABC, ACB, BAC, BCA, CAB, and CBA result in equivalent stack characteristics.
@@
Figure 13: (a) Showing each stacking sequence for the two identical samples with total deposited ABC stack thickness of 60 nm and thickness of the fused silica substrate is 1 mm. (b) Showing angle-dependent measurements for three distinct stacking sequences of the ABC sample, both for p-polarized fundamental and second harmonic. Depending on the stacking sequence configuration, the SH response varies significantly. Both $F$ with an average pump power $P=0.5$ W and SH are p-polarized.
4 Conclusion